# The NEMESIS Catalogue of Young Stellar Objects for the Orion Star Formation Complex

### v1

@@ 5.2.4 Further contamination
As also suggested by our comparison with Simbad, Fig. 10 also reveals contamination by low-mass MS stars. We isolated this contamination based on a 20 Myr MESA isochrone, finding 2,845 sources below the isochrone. Based on the amount of Gaia DR3 good-quality data available, this suggests a % contamination incidence.
5.3 Gaia’s RUWE in the context of a large sample of YSOs
As an illustrative application of our catalogue, we further contribute to previous suggestions that Gaia’s RUWE may be inflated in YSOs. As discussed in Sec. 4.3.3, Gaia’s astrometric solution assumes that sources are point-like with a well-defined centre of light and a single-star movement. Along with multiple systems, extended sources also deviate from these assumptions due to their non-point-like nature and the underlying larger uncertainty in resolving their centre of light. In YSOs, the diffuse emission of the envelope in Class 0/I sources, or even the radiation re-emitted by a thick disk in Class II sources could be the culprit of an uncertain centre of light. In Sec. 5.3.1, we address the influence of the disk in Gaia’s DR3 RUWE.
Beyond extended sources, photometric variability has also been posed as a contributor to RUWE. For example, Belokurov et al. (2020) previously uncovered a trend of inflated RUWE as a function of variability amplitude for RRLyrae and Cepheids observed with Gaia DR2. As further discussed by these authors, the inflation of RUWE by variability can be traced back to RUWE’s definition (Lindegren, 2018, eq. (4)). As its name says, RUWE is a normalised version of Gaia’s 5-parameter astrometric solution’s unit error, where a normalisation coefficient is derived from one of the modes UWE’s as a function of magnitude and colour. In variable stars where sharp changes of magnitude and colour take place, this normalisation may be incorrect, yielding spurious inflation of RUWE. Since variability is a prevalent characteristic of YSOs (Sec. 3.8), Sec. 5.3.2 we also address the influence of variability on Gaia’s DR3 RUWE.
5.3.1 Does the presence of a disk inflate Gaia’s RUWE?
The inflation of RUWE values due to the presence of a circumstellar disk has been previously suggested by Fitton et al. (2022), based on a modest sample of 122 single stars selected from adaptative optics surveys for which they find an excess of disk-bearing sources in the RUWE range 1–2.5. Here, we contribute to this evidence by examining the RUWE of a much larger sample of YSOs. We selected such a sample based on sources from our with had both derived in Sec. 3.3.1 and Gaia’s DR3 RUWE. To maximise the sample’s purity, we removed all sources flagged as possible contaminants in Sec. 5.2. Next, we split this sample at to define a diskless sample (Class III sources) and a disk-bearing sample (Class II, FS and Class 0/I sources). We note that, unfortunately, the requirement for Gaia data will inexorably remove the least evolved and more embedded Class I/O sources in our catalogue, as those are too extinct and beyond Gaia’s detection capacity.
Next, we also removed all sources flagged as multiple in Sec. 4, except for cases where the only multiplicity flag was the RUWE-based one, R (Sec. 4.3.3). This is required to preserve sources in the RUWE range that Fitton et al. (2022) claim to be populated by disk-bearing sources. Although 51% of the R labelled sample is still removed, as they were also identified as multiple systems by other methods, a modest fraction of sources with RUWE and probably real unresolved systems are left in the sample. At this point, we noticed that this fraction was larger for the disk-bearing () than for the diskless () sample, which is likely a result of the other methods of detection of unresolved pairs being more efficient in detecting binaries around diskless stars (for example, eclipsing binaries are easier to identify when the erratic variability of disk-bearing YSOs’ is not present in the light curve). To further minimise the influence of remaining binaries in our results, we focused our analysis on the RUWE2.5 range, which has been suggested to be the relevant range for disk-bearing sources (see Fitton et al., 2022). Our final sample was composed of 8,962 diskless sources and 1,593 disk-bearing sources.
Finally, we employed a series of non-parametric statistical tests (Kolmogorov-Smirnov, Mann-Whitney U, and Anderson-Darling) to examine the differences between RUWE distributions of disk-bearing and diskless sources, finding unanimous support to reject the null hypothesis that the two samples originate from the sample parent distribution. To further support this evidence, we have also repeated the investigation for samples of disk-bearing and diskless sources defined as in Fitton et al. (2022) where an IR-excess colour criterion based on our clean ALLWISE data (as of App. A, , is used. We have also considered more restrictive selection criteria, where we use both the and the AllWISE colour selection criteria to select disk-bearing sources and where we use for selecting diskless sources without reminiscent thin disks. In all cases, statistical tests support the difference in the RUWE distributions for disk-bearing and diskless sources. These differences are also illustrated in Fig. 12, where an excess of sources with larger RUWE is clearly visible in the normalised and cumulative distributions of RUWE. Moreover, the left panel of Fig. 12 shows the distribution of RUWE as a function of with the rolling interquartile range of the distribution indicating a more robust tail toward large RUWE values for more positive values of (larger contribution of the disk to the IR-SED).
5.3.2 Does the variability of YSOs inflate RUWE?
Next, we use Gaia DR3 calculated in Sec. 3.8.1 (available for of the sources in our catalogue) to add to the evidence for variability influence on RUWE (e.g. Belokurov et al., 2020; Barber & Mann, 2023). As in the previous section, we minimised the incidence of binaries in the sample by removing all sources from Sec. 4 with a multiplicity flag other than R and examining RUWE values up to 2.5. We also removed all sources flagged as possible contaminants in Sec. 5.2. which left us with 10,326 sources. Next, we bin the data at magnitude ranges (848 sources), (3,571), (3,850), and : (2,057) with the first two bins following Gaia’s window classes, and the remaining aimed at minimising the effects of the signal-to-noise variations with magnitude. For each magnitude bin, we estimate the first, second, and third rolling quartiles for the distributions of RUWE as functions of . The results are shown in Fig. 13, where a variation of in RUWE as a function of amplitude is verified.
6 Summary
@@
We have assembled the largest historical compilation of YSOs in the Orion Star Formation Complex to date, including 28,879 previously reported YSO candidates. The list of YSO candidates reported here included:
@@
78% of sources included spectroscopic features associated with the star-disk-interaction and accretion processes in YSOs;
@@
compiling SEDs for all sources in the catalogue with 96% of sources in the catalogue with at least 5 data points in their SEDs;
@@
homogeneously deriving indices and IR classes for 93% of sources in the catalogue;
@@
extensively evaluating the multiplicity of sources, resulting in 43% sources in the catalogue labelled as likely multiple sources (binary candidates or visual pairs).
Due to the positive evidence approach adopted for our data collection, our catalogue prioritises completeness over purity. We estimate that of the sources in our catalogue are reliable YSOs, with 0.3% of the catalogue is composed of massive stars, 11% are likely giants contaminants, are likely extragalactic contaminants, and 12.5% are likely MS contaminants. Finally, we have estimated that the incidence of duplicates in this catalogue is below the 1% level.
As an illustrative application of the catalogue, we add to the evidence that Gaia’s RUWE parameter - commonly used for identifying unresolved binaries - may be inflated in YSOs due to both the presence of a disk and their large variability.
The present catalogue is already in use by our collaborators for upcoming publications related to the NEMESIS project. (Hernandez et al., 2024 submitted) employed the catalogue to select a list of YSOs to investigate the morphology of YSOs using self-organising maps; (Marton et al., 2024b, 2025 in prep) used the catalogue as a training set for the identification of YSOs all-sky using deep learning techniques; and (Mas et al., 2025 in prep) used the YSO observables collated here to validate the results concerning the variability of YSOs observed by Gaia DR3. Additionally, two ongoing studies will add value to the present catalogue through the fitting of YSO’s spectral energy distributions (SEDs) to synthetic libraries based on radiative transfer models (Gezer et al., 2025 in prep), and the ranking of bona fide YSOs (Roquette et al. in prep. 2025).
@@ Acknowledgements.
- Gezer et al. (2025 in prep) Gezer, I., Roquette, J., Marton, G., et al. 2025 in prep, A&A
@@
- Madarász et al. (2024 submitted) Madarász, M., Marton, G., Gezer, I., et al. 2024 submitted, A&A
@@
- Mas et al. (2025 in prep) Mas, C., Roquette, J., Audard, M., et al. 2025 in prep, A&A
@@ Appendix A ALLWISE Cleaning
Our RF classifiers888We experimented with various machine learning classification methods, including also Logistic Regression, k-Neighbours, Naive Bayes, Decision Tree, Support Vector Machine, and Neural Networks, but found that the RF classifier yields the highest classification accuracy during the testing phase. were trained on samples of 3,800 labelled sources (with 50-50% labelled as true and false detections), using 21 features per band: AllWISE profile-fit magnitudes and their uncertainties (w?mpro, w?sigmpro), signal-to-noise (w?snr) and profile fit reduced chi-squared (w?chi2), the magnitudes and uncertainties at 8 different apertures (w?mag_# and w?sigm_#, where # refer to the 8 available apertures within AllWISE), and the ratios between magnitudes at consecutive apertures. Classification reports for both bands are given in Table 2 and Confusion matrices are shown in Fig. 14.
@@
Appendix B Summary of studies included in the historical compilation
@@
| Hillenbrand et al. (1998) | EWacc disk, SED, | x | ||||
@@
| Ali & Noriega-Crespo (2004) | disk, SED | x | x | |||
@@
| Hernández et al. (2007b) | disk, SED | x | ||||
| Barrado y Navascués et al. (2007) | disk, SED | x | ||||
| Hernández et al. (2007a) | disk, SED | x | x | |||
@@
| Fang et al. (2009) | Li, SpT, EWacc, H, disk, SED, | x | x | |||
@@
| Mookerjea & Sandell (2009) | disk, SED, | x | ||||
| Hernández et al. (2009) | disk, SED, | x | ||||
@@
| Connelley & Greene (2010) | SpT, EWacc, veiling, bin, disk, SED, | x | ||||
@@
| Rigliaco et al. (2011) | , H, disk, SED, | x | x | |||
| Barrado et al. (2011) | X-ray, disk, SED, | x | ||||
| Franciosini & Sacco (2011) | X-ray, SpT, disk, | x | ||||
@@
| Caballero et al. (2012) | Li, SpT, EWacc, H, disk, | x | x | |||
| Bayo et al. (2012) | rotation, SpT, EWacc, H, disk, | x | x | |||
@@
| Kryukova et al. (2012) | disk, SED, | x | ||||
| Megeath et al. (2012) | disk, SED, | x | ||||
@@
| Hsu et al. (2013) | Li, RV, SpT, disk, SED, | x | ||||
| Stutz et al. (2013) | , bin, disk, SED, | x | ||||
@@
| Kim et al. (2013) | X-ray, , SpT, bin, disk, | x | ||||
| Fang et al. (2013) | Li, SpT, EWacc, H, disk, SED, | x | ||||
@@
| Hernández et al. (2014) | Li, RV, SpT, EWgravity, H, bin, disk, SED, | x | x | x | ||
@@
| Spezzi et al. (2015) | disk, SED, | x | ||||
| Koenig et al. (2015) | SpT, EWacc, H, disk, SED, | x | ||||
@@
| Hasenberger et al. (2016) | SpT, disk, | x | ||||
| Megeath et al. (2016) | disk, SED, | x | ||||
@@
| Kounkel et al. (2016b) | bin, disk, | x | ||||
| Lewis & Lada (2016) | bin, disk, | x | ||||
| Furlan et al. (2016) | , disk, SED, | x | ||||
| Kim et al. (2016) | Li, , SpT, H, bin, disk, | x | ||||
@@
| Fang et al. (2017) | Li, SpT, H, disk, | x | ||||
| Suárez et al. (2017) | Li, , SpT, H, disk, | x | ||||
@@
| Grant et al. (2018) | SpT, disk, SED, | x | ||||
| Yao et al. (2018) | , logg, EWacc, disk, SED, | x | ||||
| Cottle et al. (2018) | disk, SED, | x | ||||
@@
| Kounkel et al. (2019) | RV, , logg, veiling, bin, disk, | x | ||||
@@
| Pinzón et al. (2021) | rotation, RV, , SpT, bin, disk, | x | x | |||
@@
| Smith et al. (2023) | rotation, , disk, | x |
Appendix C Database Summary
In this appendix, we outline the data types included in our compilation and summarise the various tables that compose our catalogue.
@@ Appendix C Database Summary
| Internal_ID | NEMESIS Internal ID | - | - | - |
| Simbad | ||||
| main_id | Main identifier in Simbad | - | 28,233 | 25,530 |
| otype | Main Object Type in Simbad | - | - | - |
| otypes | Full list of Object types in Simbad | - | - | - |
| ra_simbad, dec_simbad | RA, Dec coordinates in Simbad | degree | - | - |
@@
| sep2simbad | Separation to Simbad counterpart | arcsec | - | - |
| Contamination | ||||
| massive_star | Likely massive star (Sec. 5.2.1) | - | 80 | 80 |
| ms_contaminant | Likely MS contaminant(Sec. 5.2.4) | - | 2,845 | 2,845 |
| giant_label | Labelled Giant contaminant (App. D.2.1 and Sec. 5.2.2) | - | 457 | 457 |
| giant_prob | Probability of source being a Giant (App. D) | - | 20,714 | 20,714 |
| extraglactic_label | Labelled extragalactic contaminant (Sec. 5.2.3 and App. D.2.1) | - | 164 | 164 |
| extraglactic_prob | Probability of source being extragalactic (App. D) | - | 16,079 | 16,079 |
@@
| Temperature | ||||
| Teff | Effective temperature | K | 71,135 | 17,443 |
@@
| Standard Classification Scheme | ||||
| alpha_2_24 | estimated in Sec. 3.3.1. | - | 27,799 | 27,799 |
@@
| Literature values | ||||
| alpha | compiled from the literature | - | 48,657 | 5,379 |
| class | IR class compiled from the literature | - | 16,716 | 8,097 |
| IR_range | wavelength range used for classification | - | ||
@@
| disk | Disk description keyword | - | 8,054 | 4,096 |
@@
| logg | derived from spectra fit | dex | 57,649 | 12,313 |
@@
| H | ||||
| EW_Ha | EW of at Å | Å | 32,760 | 20,667 |
| FW10%(H) | Full-width at 10 of the H emission profile peak | km/s | 1,181 | 945 |
@@
| veiling_r_h | ir-veiling at the H-band (m) | - | 32,132 | 9,031 |
@@
| veiling_r_r7465 | veiling around 7465Å | - | 1,444 | 361 |
@@
| Variability | ||||
| Per | Variability periods | day | 21,932 | 10,147 |
@@
| AG_proxy | Gaia DR3 -proxy variability index | - | 24,632 | 24,632 |
| Rotation | ||||
| vsini | km/s | 20,820 | 10,651 | |
@@
| Fx | X-ray flux | erg/cm2/s | 2,371 | 2,333 |
| logLx | Total X-ray luminosity | erg/s | 6,121 | 3,051 |
@@
| logNH | Hydrogen column density | 4,324 | 1,952 | |
| HR# | Hardness ratio #, # in (1, 2, or 3) | - | 1,364 | 1,364 |
@@
Appendix D Extragalactic and Giants contamination assessment

### v2

@@ 5.2.4 Further contamination
As also suggested by our comparison with Simbad, Fig. 12 also reveals contamination by low-mass MS stars. We isolated this contamination based on a 20 Myr MESA isochrone, finding 2 845 sources below the isochrone. Based on the amount of Gaia DR3 good-quality data available, this suggests a % contamination incidence.
6 Summary of Results
@@
We have assembled the largest historical compilation of YSOs in the Orion Star Formation Complex to date, including 27 879 previously reported YSO candidates. The list of YSO candidates reported here included:
@@
78% of sources included spectroscopic features associated with the star-disc-interaction and accretion processes in YSOs;
@@
compiling SEDs for all sources in the catalogue, with 96% of sources with at least 5 data points in their SEDs;
@@
homogeneously deriving indices and IR classes for 93% of sources;
@@
extensively evaluating the multiplicity, resulting in 43% sources in the catalogue labelled as likely multiple sources (binary candidates or visual pairs).
Due to the positive evidence approach adopted for our data collection, our catalogue prioritises completeness over purity. We estimate that of the sources in our catalogue are reliable YSOs, with 0.3% of the catalogue is composed of massive stars, 13% are likely giants contaminants, are likely extragalactic contaminants, and 12.5% are likely MS contaminants. Finally, we have estimated that the incidence of duplicates in this catalogue is below the 1% level.
As an illustrative application of the catalogue, we add to the evidence that Gaia’s RUWE parameter - commonly used for identifying unresolved binaries - may be inflated in YSOs due to their large variability.
The present catalogue is already in use by our collaborators for upcoming publications related to the NEMESIS project. (Hernandez et al., 2024 submitted) employed the catalogue to select a list of YSOs to investigate the morphology of YSOs using self-organising maps; (Marton et al., 2024b, 2025 in prep) used the catalogue as a training set for the identification of YSOs all-sky using deep learning techniques; and (Mas et al., 2025) used the YSO observables collated here to validate the results concerning the variability of YSOs observed by Gaia DR3. (Gezer et al., 2025) presented further YSO parameters by fitting YSOs’ SEDs with disc models. An ongoing study will further add value to the present catalogue through the ranking of bona fide YSOs (Roquette et al. in preparation).
@@ Acknowledgements.
- Gezer et al. (2025) Gezer, I., Marton, G., Roquette, J., et al. 2025, A&A, 696, A196
@@
- Madarász et al. (2025) Madarász, M., Marton, G., Gezer, I., et al. 2025, A&A, 696, A37
@@
- Mas et al. (2025) Mas, C., Roquette, J., Audard, M., et al. 2025, arXiv e-prints, arXiv:2503.23530
@@ Appendix A ALLWISE Cleaning
Our RF classifiers999We experimented with various machine learning classification methods, including also Logistic Regression, k-Neighbours, Naive Bayes, Decision Tree, Support Vector Machine, and Neural Networks, but found that the RF classifier yields the highest classification accuracy during the testing phase. were trained on samples of 3,800 labelled sources (with 50-50% labelled as true and false detections), using 21 features per band: AllWISE profile-fit magnitudes and their uncertainties (w?mpro, w?sigmpro), signal-to-noise (w?snr) and profile fit reduced chi-squared (w?chi2), the magnitudes and uncertainties at 8 different apertures (w?mag_# and w?sigm_#, where # refer to the 8 available apertures within AllWISE), and the ratios between magnitudes at consecutive apertures. Classification reports for both bands are given in Table 2 and Confusion matrices are shown in Fig. 14.
@@
Appendix B reliability flags
To quantify the effect of the reduced availability of mid/far infrared data on our -indices (Sect. 3.3.1), we carried out an undersampling exercise. Recalling our requirement of two data points at least 2 away in wavelength for derivation, we selected a testing sample of 16 282 sources that followed this requirement for wavelengths m, and had at least one additional data point in the range m. For this sample, we derived an additional -index using only the data between m. Fig. 16 shows a comparison of this index to our main . For the specific discussion in this Appendix, we limit our classes to three: discless, () thin disc () and thick disc (). We find that of all sources had the same class with either definition (white and gray-shaded areas in Fig. 16); sources were discless sources miss-detected as thin disc (yellow); were thick discs miss-detected as thin disc (green); were thin discs miss-detected as thick disc (red); of sources were thin discs miss-detected as discless (purple); were thick discs miss-detected as discless (blue); and sources were discless miss-detected as thick discs (cyan). With 33% of our YSO candidates having no SED datapoints for m, we expected such bias to affect of our -indices. We note that this estimation does not take into account domain-motivated choices of the wavelength range used for , and we point the interested reader to Großschedl et al. (2019) for an in-depth discussion of the effect of definitions on disc classification. We also note that the numbers discussed so far refer to the total number of sources analysed. In the bottom plot of Fig. 16, we detailed miss-classification rates normalised by the number of sources in each class derived with . Although only 71.2 and 66.9% of the original thick discs and diskless sources are recovered as such, these classes have relatively large purity, with 93.7% of thick disks and 97.8% of discless sources correctly detected as such with the index. In contrast, while 88.6% of thin discs are recovered as such, this class has a much smaller purity of 74.8%.
Appendix C Summary of studies included in the historical compilation
@@
| Hillenbrand et al. (1998) | EWacc disc, SED, | x | ||||
@@
| Ali & Noriega-Crespo (2004) | disc, SED | x | x | |||
@@
| Hernández et al. (2007b) | disc, SED | x | ||||
| Barrado y Navascués et al. (2007) | disc, SED | x | ||||
| Hernández et al. (2007a) | disc, SED | x | x | |||
@@
| Fang et al. (2009) | Li, SpT, EWacc, H, disc, SED, | x | x | |||
@@
| Mookerjea & Sandell (2009) | disc, SED, | x | ||||
| Hernández et al. (2009) | disc, SED, | x | ||||
@@
| Connelley & Greene (2010) | SpT, EWacc, veiling, bin, disc, SED, | x | ||||
@@
| Rigliaco et al. (2011) | , H, disc, SED, | x | x | |||
| Barrado et al. (2011) | X-ray, disc, SED, | x | ||||
| Franciosini & Sacco (2011) | X-ray, SpT, disc, | x | ||||
@@
| Caballero et al. (2012) | Li, SpT, EWacc, H, disc, | x | x | |||
| Bayo et al. (2012) | rotation, SpT, EWacc, H, disc, | x | x | |||
@@
| Kryukova et al. (2012) | disc, SED, | x | ||||
| Megeath et al. (2012) | disc, SED, | x | ||||
@@
| Hsu et al. (2013) | Li, RV, SpT, disc, SED, | x | ||||
| Stutz et al. (2013) | , bin, disc, SED, | x | ||||
@@
| Kim et al. (2013) | X-ray, , SpT, bin, disc, | x | ||||
| Fang et al. (2013) | Li, SpT, EWacc, H, disc, SED, | x | ||||
@@
| Hernández et al. (2014) | Li, RV, SpT, EWgravity, H, bin, disc, SED, | x | x | x | ||
@@
| Spezzi et al. (2015) | disc, SED, | x | ||||
| Koenig et al. (2015) | SpT, EWacc, H, disc, SED, | x | ||||
@@
| Hasenberger et al. (2016) | SpT, disc, | x | ||||
| Megeath et al. (2016) | disc, SED, | x | ||||
@@
| Kounkel et al. (2016b) | bin, disc, | x | ||||
| Lewis & Lada (2016) | bin, disc, | x | ||||
| Furlan et al. (2016) | , disc, SED, | x | ||||
| Kim et al. (2016) | Li, , SpT, H, bin, disc, | x | ||||
@@
| Fang et al. (2017) | Li, SpT, H, disc, | x | ||||
| Suárez et al. (2017) | Li, , SpT, H, disc, | x | ||||
@@
| Grant et al. (2018) | SpT, disc, SED, | x | ||||
| Yao et al. (2018) | , logg, EWacc, disc, SED, | x | ||||
| Cottle et al. (2018) | disc, SED, | x | ||||
@@
| Kounkel et al. (2019) | RV, , logg, veiling, bin, disc, | x | ||||
@@
| Pinzón et al. (2021) | rotation, RV, , SpT, bin, disc, | x | x | |||
@@
| Smith et al. (2023) | rotation, , disc, | x |
Appendix D Database Summary
The database curated as part of the NEMESIS Catalogue of YSOs for the OSFC can be accessed via SQL (at https://www.astro.unige.ch/nemesis/) and through CDS. The 19 thematic tables with data compiled in the present study are organised in 12 directories. A full description of directories, sub-tables, column names, units and data availability is also available as part of the online material (nemesis_osfc_description.csv). This appendix further summarises key data products, the number of individual measurements collected for each observable, and the number of sources with at least one measurement. We stress that the data is organised following a relational database paradigm. Each source is associated with a unique NEMESIS_ID key. This unique source identifier should be used to group multiple measurements (when available) for a given source and match different thematic tables.
@@ Appendix D Database Summary
Tab. ir/ : Quantities related to IR YSO classification schemes discussed Sect. 3.3;
standard_classification_nemesis.csv YSO classification derived in this study.
disk_literature.csv: IR-based YSO classification collected from the literature.
Tab. : accretion/ : Quantities related to accretion or material inflow/outflow (Sect. 3.6);
Tab. : quantities related to the X-ray emission of YSOs (Sect. 3.9);
Tab. multiplicity/multiplicity_label.csv: Multiplicity labels (Sect. 4).
@@
| Internal_ID | NEMESIS Catalogue Internal ID | - | 27 879 | 27 879 |
| Simbad counterparts (main/simbad.csv) | ||||
| main_id | Main identifier in Simbad | - | 33 846 | 25 902 |
| otype | Main Object Type in Simbad | - | 33 690 | 25 901 |
| otypes | Full list of Object types in Simbad | - | 33 615 | 25 902 |
| ra_simbad, dec_simbad | RA, Dec coordinates in Simbad | degree | 33 846 | 25 902 |
@@
| counterpart_groupID | identifier for grouping multiple possible counterparts | - | 33 846 | 25 902 |
| sep2simbad | Separation to Simbad counterpart | arcsec | 12 835 | 4 891 |
| Contamination (main/contamination.csv)) | ||||
| massive_star | Flag indicative of likely massive star (Sect. 5.2.1) | - | 80 | 80 |
| ms_contaminant | Likely MS contaminant(Sect. 5.2.4) | - | 2 845 | 2 845 |
| extraglactic_prob | Probability of source being extragalactic (App. E) | - | 15 855 | 15 855 |
| extraglactic_pred | Extragalactic predicted label (App. E) | - | 15 855 | 15 855 |
| extraglactic_label | Labelled extragalactic contaminant (Sect. 5.2.3 and App. E.2.1) | - | 125 | 125 |
| giant_prob | Probability of source being a Giant (App. E) | - | 20 546 | 20 546 |
| giant_prob | Probability of source being a Giant (App. E) | - | 20 546 | 20 546 |
| giant_label | Labelled Giant contaminant (App. E.2.1 and Sect. 5.2.2) | - | 490 | 490 |
[…]
@@
| Temperature (hrd/temperature.csv) | ||||
| Teff | Effective temperature | K | 73 808 | 17 589 |
@@
| field | description | units |
[…]
| nu_Fnu | observed flux density | |
| nu_Fnu_error | flux density uncertainty | |
| wavelength | effective wavelength | Å |
| flux | observed flux | Jy |
| freq | frequency | Hz |
| bib | NASA/ADS bibcode for the original paper linked to the reported quantity | - |
| filter | SVO filter name | - |
| SED summary table (sed/sed_summary.csv) | ||
| uv | Number of SED datapoints in the range Å | - |
| optical_1 | Number of SED datapoints in the range Å | - |
| optical_2 | Number of SED datapoints in the range Å | - |
| optical_3 | Number of SED datapoints in the range Å | - |
| nir_1 | Number of SED datapoints in the range Å | - |
| nir_2 | Number of SED datapoints in the range Å | - |
| nir_3 | Number of SED datapoints in the range Å | - |
| mir_1 | Number of SED datapoints in the range Å | - |
| mir_2 | Number of SED datapoints in the range Å | - |
| mir_3a | Number of SED datapoints in the range Å | - |
| mir_3b | Number of SED datapoints in the range Å | - |
| mir_4 | Number of SED datapoints in the range Å | - |
| mir_5 | Number of SED datapoints in the range Å | - |
| fir_1 | Number of SED datapoints in the range Å | - |
| fir_2 | Number of SED datapoints in the range Å | - |
| fir_3 | Number of SED datapoints in the range Å | - |
@@
| Standard Classification Scheme (ir/stnadard_classification_nemesis.csv) | ||||
| alpha_2_24 | estimated in Sect. 3.3.1. | - | 25 799 | 25 799 |
@@
| lambda_min_2_24 | Minimum wavelength available for estimation | m | 25 799 | 25 799 |
| lambda_max_2_24 | Maximum wavelength available for estimation | m | 25 799 | 25 799 |
| Literature values (disk_literature.csv) | ||||
| alpha | compiled from the literature | - | 48 657 | 5 379 |
| class | IR class compiled from the literature | - | 16 716 | 8 097 |
| IR_range | wavelength range used for classification | - | 62 986 | 9 438 |
@@
| disc | disc description keyword | - | 8 054 | 4 096 |
@@
| (gravity/logg_DB.csv) | ||||
| logg | derived from spectra fit | dex | 57 649 | 12 313 |
| line measurements (gravity/ew_gravity.csv) | ||||
@@
| H (accretion/Ha_DB.csv) | ||||
| EW_Ha | EW of at Å | Å | 32 760 | 20 667 |
| FW10%(H) | Full-width at 10 of the H emission profile peak | km/s | 1 181 | 945 |
| EW_Ha_PHOT | EW of from photometry | Å | 1 514 | 1 239 |
[…]
| other lines(accretion/EW_accretion_other_DB.csv) | ||||
@@
| veiling_r_h | ir-veiling at the H-band (m) | - | 32 132 | 9 031 |
@@
| veiling_r_r7465 | veiling around 7465Å | - | 1 444 | 361 |
@@
| Variability (variability/amplitude.csv) | ||||
| AG_proxy | Gaia DR3 -proxy variability index | - | 24 632 | 24 632 |
| key | description | units | count | # sources |
| Rotation (rotation/rotation.csv) | ||||
| Per | Variability periods | day | 21 932 | 10 147 |
@@
| vsini | km/s | 20 820 | 10 651 | |
@@
| xray/Xray_DB.csv | ||||
| Fx | X-ray flux | erg/cm2/s | 2 371 | 2 333 |
| logLx | Total X-ray luminosity | erg/s | 6 121 | 3 051 |
@@
| logNH | Hydrogen column density | 4 324 | 1 952 | |
| HR# | Hardness ratio #, # in (1, 2, or 3) | - | 1 364 | 1 364 |
@@
| key | description | unit | count | # sources |
| multiplicity/multiplicity_label.csv | ||||
[…]
| bin_ext_type | Multiplicity type in source paper | - | 18 659 | 12 102 |
Appendix E Extragalactic and Giants contamination assessment